OpenStudy (hexagon001):

how to find the Taylor series of the following: f(z)=z^2 + zcosz [in the context of complex analysis] I am getting the following...

OpenStudy (hexagon001):

\[z^2+z(1-\frac{ z^2 }{ 2! }+\frac{ z^4 }{ 4! }-\frac{ z^6 }{ 6! }-...)\]

OpenStudy (agent0smith):

I'm guessing it's a maclaurin series (about x=0)...?

OpenStudy (hexagon001):

Never heard of maclaurin series...... i should mention that my question is in the context of complex analysis

OpenStudy (agent0smith):

Ah, I don't think I've encountered complex analysis. This may help, though... http://math.fullerton.edu/mathews/c2003/TaylorSeriesMod.html

OpenStudy (hexagon001):

Thanks. I checked on wolfram and i got the correct series. http://www.wolframalpha.com/input/?i=taylor+series+%28z%5E2+%2B+z+cosz+%29 thanks again.

OpenStudy (agent0smith):

Hmm i guess the complex taylor series is the same as the non complex one. Btw a maclaurin series is just a taylor series centered at x=0 (or z=0 in this case).

OpenStudy (hexagon001):

oh yes this series was about 0 but the name Maclaurin wasnt used..but good to know. thanks